### v1

# Spontaneous Symmetry Breaking in Graviweak Theory

Abstract
Graviweak theory seeks to unify gravity (specifically in its self-dual formulation) with the weak interaction, preying on their parallel chiral $SU(2)$ structures. In this paper we further this idea by folding it with the concept of spontaneous symmetry breaking. We do this first with a standard Higgs field and potential, starting with a unifying parity-invariant theory which splits into the usual gravity and weak sector under spontaneous symmetry breaking. By rewriting the theory in the two-spin framework we are then prompted to discuss generalizations, within the generic approach known as MacDowell-Mansouri theories where a larger internal gauge group is broken. One of the predictions of the ensuing construction is a non-minimal coupling in the low energy broken phase between curvature and the weak gauge fields, translating at the quantum level to a direct channel between the graviton and the weak bosons.
I Introduction and Motivation
The unification of gravity with the other fundamental interactions remains one of the most profound challenges in theoretical physics. While the Standard Model successfully unifies the electroweak and strong interactions within a gauge-theoretic framework, gravity remains conceptually and technically distinct due to its geometric nature. An intriguing approach to unification is the idea that gravity and the weak interaction could emerge from a common underlying chiral gauge theory. This framework is referred to as graviweak unification Alexander:2012ge ; Alexander:0706.4481 ; Nesti:0706.3307 ; Nesti:0706.3304 . The motivation stems from the realization that both gravity and the weak interaction can be formulated as chiral theories. The weak interaction is intrinsically chiral—only left-handed fermions couple to the weak gauge bosons—while gravity, in the Einstein-Cartan formalism, can also be made chiral by splitting the spin connection into left- and right-handed components associated with the group structure KirillBook ; Freidel:2005mq ; Capovilla:1991qt ; Plebanski:1977pw ; Celada:2016jdt ; Alexandre:2407.19363
$SO(4)\cong SU(2)_{L}\times SU(2)_{R}.$
The parity asymmetry of the weak interaction suggests that parity should also be broken in the gravitational sector at a high energy scale. If the chiral asymmetry of the weak interaction arises from spontaneous symmetry breaking Wilczek:1998ea , a similar mechanism could break parity in the gravitational sector, leaving only a left-chiral gravitational connection at low energies.
These considerations motivate the construction of a theory in which gravity is initially described by both chiral components. A spontaneous symmetry-breaking mechanism then suppresses the right-handed chiral component at low energies. The surviving left-handed component would naturally couple to left-handed fermions, mimicking the chiral nature of the weak interaction. In particular, the gravitational interaction should have the opposite chirality of the weak interaction, reflecting the parity-breaking structure of the Standard Model. This paper develops a model of graviweak unification where a Higgs-like field higgs generates this parity asymmetry dynamically through a $Z_{2}$ symmetric potential. The right-handed kinetic term becomes heavy and decouples, leaving a chiral gravitational sector.
We develop these ideas in this paper along the following logical thread. In Section II we propose a gravitational theory formulated in terms of chiral components of the Lorentz connection and tetrad fields. The total Lagrangian includes Einstein-Hilbert-type terms, Yang-Mills-like kinetic terms for the chiral curvatures, and a Higgs-like antisymmetric tensor field $\phi^{ab}$ that couples to the chiral sectors and dynamically breaks parity symmetry through its vacuum expectation value. Thus, we have a theory that unifies gravity and the weak interaction in a parity invariant construction before symmetry breaking, reducing to the Ashtekar self-dual formulation of gravity plus the weak sector after symmetry breaking. This will be the foundations of this paper, around which we will investigate developments and variations.
In Section III we then rewrite this theory in the two-spinor notation, the natural natural language for treating gravity and chiral gauge fields on the same footing. This formalism is especially convenient for formulating unification frameworks such as graviweak models and it points us towards a new development: parity symmetry breaking that goes hand in hand with a larger gauge group. This receives further exploration in Section IV, where we extend the internal Lorentz group underpinning the basic construction to $SO(1,3+N)$. A major implication is the appearance of possible interactions between the gravitational 2-form curvature $R^{a}_{~{}b}$ of $SO(1,3)$ and the Yang-Mills field strength $F^{ij}$ of a $SO(N)$ group. This is a remarkable novelty with regards to previous graviweak models. We close this paper with a few speculations on the testability and falsifiability of this scenario.
II The Graviweak-Higgs System
We start by laying down the structure of the theory. The theory is initially parity invariant and built from left- and right-handed chiral curvature 2-forms, defined as:
$R^{(\pm)}_{ab}=d\omega^{(\pm)}_{ab}+\omega^{(\pm)}_{a}{}^{c}\wedge\omega^{(\pm)}_{cb}$ (1)
where the chiral components of the Cartan spin connection $\omega_{ab}$ are given by:
$\omega^{(\pm)}_{ab}=\frac{1}{2}\left(\omega_{ab}\pm\frac{i}{2}\epsilon_{ab}{}^{cd}\omega_{cd}\right).$ (2)
We then consider the most general parity-invariant gauge theory containing only linear and quadratic terms in the curvature, with the condition that the theory should result in general relativity and the weak sector after spontaneous symmetry breaking. Its Lagrangian is:
$\mathcal{L}_{\text{total}}=\;$ $\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge\left(R^{(+)}_{cd}+R^{(-)}_{cd}\right)$
$+\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4g_{-}^{2}}R^{(-)}_{ab}\wedge\star R^{ab(-)}$
$+\frac{\alpha}{2}\,\phi^{bc}\left(R^{(+)}_{ab}\wedge\star R^{a(+)}_{~{}c}+R^{(-)}_{ab}\wedge\star R^{a(-)}_{~{}c}\right)$
$-\frac{1}{2}D\phi^{ab}\wedge*D\phi_{ab}+\frac{V(\phi)}{24}\epsilon^{abcd}e_{a}\wedge e_{b}\wedge e_{c}\wedge e_{d}$ (3)
where the star symbol $\star$ denotes the Hodge dual on the frame indices and the asterisk * denotes the Hodge dual on the spacetime indices:
$*R^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}R^{ab\rho\sigma}$ (4)
and
$\star R^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon^{abcd}R_{cd\mu\nu}.$ (5)
A few comments on this action are in order. First note that the parity transformations under which this theory is initially symmetric are defined as follows. The tetrad transforms as $e^{a}(x^{0},\vec{x})\mapsto e^{a}(x^{0},-\vec{x})$, and the chiral components of the connection exchange under parity: $\omega^{(+)}_{ab}(x^{0},\vec{x})\mapsto\omega^{(-)}_{ab}(x^{0},-\vec{x})$, and vice versa. Consequently, the curvatures transform as $R^{(+)}_{ab}(x^{0},\vec{x})\mapsto R^{(-)}_{ab}(x^{0},-\vec{x})$ and $R^{(-)}_{ab}(x^{0},\vec{x})\mapsto R^{(+)}_{ab}(x^{0},-\vec{x})$. The scalar field $\phi^{ab}(x^{0},\vec{x})$ is assumed to be parity-even, transforming as $\phi^{ab}(x^{0},\vec{x})\mapsto\phi^{ab}(x^{0},-\vec{x})$.
Next note that the Einstein-Hilbert-like term is constructed as a sum over both chiral sectors and takes the form:
$\mathcal{L}_{\text{grav}}=\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge\left(R^{(+)}_{cd}+R^{(-)}_{cd}\right),$
ensuring that parity symmetry is respected at the level of the classical action. The Yang-Mills-type kinetic terms are introduced independently for each chiral component and are written as:
$\mathcal{L}_{\text{YM}}=\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4g_{-}^{2}}R^{(-)}_{ab}\wedge\star R^{ab(-)}.$
These terms preserve parity symmetry provided the chiral sectors are treated on equal footing.
To introduce the possibility of spontaneous parity violation, we couple an antisymmetric tensor scalar field $\phi^{ab}$ to the difference of the two chiral kinetic terms. This field transforms in the adjoint representation of the local Lorentz group and is assumed to be real. Its coupling to the curvature is written as
$\mathcal{L}_{\text{break}}=\frac{\alpha}{2}\,\phi^{bc}\left(R^{(+)}_{ab}\wedge\star R^{a(+)}_{~{}c}+R^{(-)}_{ab}\wedge\star R^{a(-)}_{~{}c}\right),$
where $\alpha$ is a dimensionless coupling constant and $\eta_{cd}$ is the Minkowski metric on the internal indices. This term is manifestly parity-invariant under the simultaneous exchange (+)()(+$\leftrightarrow$(-)( + ) ↔ ( - ) and the invariance of $\phi^{ab}$.
Finally, the dynamics of the scalar field is governed by a standard kinetic term and a Higgs-like potential. The kinetic term is constructed using the Lorentz-covariant derivative $D\phi^{ab}=d\phi^{ab}+\omega^{a}{}_{c}\phi^{cb}+\omega^{b}{}_{c}\phi^{ac}$, and takes the form $-\frac{1}{2}D\phi^{ab}\wedge*D\phi_{ab}$. The potential is given by
$V(\phi)=\frac{\lambda}{4}\left(\phi^{ab}\phi_{ab}-v^{2}\right)^{2},$
where $\lambda$ is a coupling constant and $v$ sets the symmetry-breaking scale. When $\phi^{ab}$ acquires a vacuum expectation value, this induces a spontaneous breaking of parity symmetry by generating unequal effective dynamics for the left- and right-handed chiral sectors.
This theory is thus parity symmetric at the level of the action, and explicitly so before spontaneous symmetric breaking. However, once the scalar field condenses and selects a preferred direction in the internal space, the coupling between $\phi$ and the chiral curvatures breaks the parity symmetry spontaneously, leading to asymmetric effective dynamics in the gravitational sector. Minimizing the potential $V(\phi)$ gives a non-zero vacuum expectation value (vev) of the form:
$\langle\phi^{ab}\rangle=\frac{v}{2}\eta^{ab}$
Substituting this into the coupling term, together with the kinetic terms yields:
$\mathcal{L}_{\text{YM}}+\mathcal{L}_{\text{break}}=\frac{1}{4}\left(\frac{1}{g_{+}^{2}}+\alpha v\right)R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4}\left(\frac{1}{g_{-}^{2}}+\alpha v\right)R^{(-)}_{ab}\wedge\star R^{ab(-)}.$ (6)
At low energies, if the symmetry-breaking scale satisfies
$v\gg\frac{1}{g_{-}^{2}}$
the coefficient of the right handed kinetic term becomes very large, effectively suppressing it. As a result, the right-handed connection becomes non-dynamical and only the left-handed kinetic term remains:
$\mathcal{L}_{\text{YM}}^{\text{eff}}=\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)},$ (7)
where we have redefined
$\frac{1}{g_{+}^{2}}+\alpha v\to\frac{1}{g_{+}^{2}}$
and so the left-handed coupling constant can be controlled by the Higgs vev.
The fermion coupling is introduced through the chiral covariant derivative:
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\star e_{a}\wedge(\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}+\bar{\psi}_{R}\gamma^{a}D^{(-)}\psi_{R}),$ (8)
where the exterior covariant derivative acting on the spinors is defined by
$D^{(\pm)}\psi_{L/R}=d\psi_{L/R}+\frac{1}{4}\omega^{(\pm)ab}\gamma_{ab}\psi_{L/R}$ (9)
where $\psi_{L/R}=P_{L/R}\psi$ are the usual left- and right-handed spinor components and $\gamma_{ab}=\gamma_{[a}\gamma_{b]}$. After symmetry breaking, the right-handed connection becomes non-dynamical, reducing the fermion coupling to the left-handed chiral component:
$\mathcal{L}^{\text{eff}}_{\text{fermion}}=\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}.$ (10)
Hence, after symmetry breaking and adding the spinor term to (3) we obtain:
$\mathcal{L}^{\text{eff}}_{\text{total}}=\;$ $\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge R^{(-)}_{cd}$
$+\frac{1}{4g_{+}^{2}}\bar{F}_{i}\wedge\star\bar{F}^{i}$
$+\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L},$ (11)
which is of course standard gravity in the Ashtekar self-dual formulation plus weak theory, with fermions. We have relabeled the weak curvature as $\bar{F}^{i}=2iR^{(+)0i}$ to put it in a more familiar form. Notice we also see that the weak sector only couples to left handed spinors, reflecting chirality and parity violation.
The equations of motion of our theory can be derived by the usual variation of the Lagrangian with respect to the fields, these being the tetrad $e^{a}$, both left- and right-handed connections $\omega^{ab(\pm)}$ and the spinor $\psi_{L}$. Hence, we get, respectively,
$\epsilon_{abcd}\left[\frac{1}{16\pi G}e^{b}\wedge R^{cd(-)}+\frac{i}{4}\bar{\psi}_{L}\gamma^{b}D^{(+)}\psi_{L}\wedge e^{c}\wedge e^{d}\right]=0$
$\star e_{a}\bar{\psi}_{L}\gamma^{a}\gamma_{cd}\psi_{L}=0$
$\epsilon_{abcd}T^{a(-)}\wedge e^{b}=0$
$\star e_{a}\gamma^{a}\wedge D^{(+)}\psi_{L}=0,$ (12)
where $T^{a(-)}=D^{(-)}e^{a}$ is the right-handed torsion of the tetrad.
In order to recover the Einstein-Cartan theory and $SU(2)$ Yang-Mills in Lorentzian signature, we will need to impose reality conditions Jacobson:1988fb . These are the following:
$\omega^{(-)}_{ab}+(\omega^{(-)}_{ab})^{*}=\omega_{ab}^{(-)}(e)$ (13)
and
$\omega^{(+)}_{ab}=(\omega^{(+)}_{ab})^{*}$ (14)
where $\omega^{(-)}_{ab}(e)$ is the right-handed spin connection which possesses vanishing torsion: $\mathrm{d}\omega^{(-)}_{ab}(e)+\omega^{(-)}_{ab}(e)\wedge e^{b}=0$.
III Graviweak Lagrangian in Two-Spinor Notation
The two-spinor notation is particularly adept for discussing parity issues so it is not surprising that it provides a natural language for treating gravity and chiral gauge fields on the same footing. This formalism is especially convenient for formulating unification frameworks such as graviweak models. As we will see towards the end of this Section, it leads to insights that justify the developments presented in the rest of this paper.
We begin by reviewing the two-spinor (or $\mathbb{C}$)) notation. In four-dimensional spacetime, the Lorentz algebra $\mathfrak{so}(3,1)$ is locally isomorphic to $\mathfrak{sl}(2,\mathbb{C})\oplus\mathfrak{sl}(2,\mathbb{C})$, corresponding to the self-dual and anti-self-dual parts of the spin connection. Using the Infeld–van der Waerden map Jacobson:1988fb ; Penrose:1984mf ; Capovilla:1991kp , spacetime indices $a=0,1,2,3$ are replaced by spinor indices via the soldering form $\sigma^{a}_{AA^{\prime}}$:
$V^{a}\longleftrightarrow V^{AA^{\prime}}=V^{a}\,\sigma_{a}^{AA^{\prime}}.$ (15)
An antisymmetric tensor such as the spin connection $\omega^{ab}$ decomposes into self-dual ()(-)( - ) and anti-self-dual (+)(+)( + ) parts as per (2), with:
$\omega^{ab}=\omega^{(+)\,ab}+\omega^{(-)\,ab}.$ (16)
These map respectively to:
$\omega^{(-)\,ab}$ $\longleftrightarrow A^{AB},$ (17)
$\omega^{(+)\,ab}$ $\longleftrightarrow\bar{A}^{A^{\prime}B^{\prime}},$ (18)
where $A^{AB}=A^{(AB)}$ is the right-chiral (SD) $\mathbb{C}$) connection, and $\bar{A}^{A^{\prime}B^{\prime}}=\bar{A}^{(A^{\prime}B^{\prime})}$ is its left-chiral (ASD) counterpart. The tetrad $e^{a}$ is similarly rewritten as:
$e^{AA^{\prime}}=e^{a}\sigma_{a}^{AA^{\prime}}.$ (19)
A detailed dictionary for the quantities appearing the last Section can then be developed and we refer the interested reader to Appendix A for details.
We can therefore translate into this language our Lagrangian (11). Using the spinor decomposition described above, we rewrite each term. The self-dual Einstein-Cartan term becomes:
$\mathcal{L}_{\text{grav}}=\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{BB^{\prime}}\wedge R^{(-)}_{AA^{\prime}BB^{\prime}},$ (20)
where $R^{(-)}_{AA^{\prime}BB^{\prime}}=\epsilon_{A^{\prime}B^{\prime}}F_{AB}$ and $F^{AB}$ is the curvature 2-form associated with the self-dual connection $A^{AB}$. Therefore we have
$\mathcal{L}_{\text{grav}}=-\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{B}_{\,\,\,A^{\prime}}\wedge F_{AB}.$ (21)
The curvature $\bar{F}^{A^{\prime}B^{\prime}}$ of the ASD connection contributes:
$\mathcal{L}_{\text{YM}}=-\frac{i}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}},$ (22)
where $\bar{F}^{A^{\prime}B^{\prime}}$ is the field strength 2-form associated with $\bar{A}^{A^{\prime}B^{\prime}}$.
The fermion kinetic term involves the covariant derivative with respect to the ASD connection and the Dirac gamma matrix $\gamma^{a}$ is translated into the soldering form:
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}},$ (23)
where $D^{(+)}$ is the spinor covariant derivative acting on left-handed spinor $\psi_{A^{\prime}}$ via the connection $\bar{A}^{A^{\prime}}_{\,\,\,\,B^{\prime}}$ and
$\star e_{AA^{\prime}}=\frac{i}{3}e_{A}^{\,\,\,\,C^{\prime}}\wedge e_{DC^{\prime}}\wedge e^{D}_{\,\,\,\,A^{\prime}}.$ (24)
The full Lagrangian in 2-spinor form becomes:
$\mathcal{L}=-\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{B}_{\,\,\,A^{\prime}}\wedge F_{AB}-\frac{i}{2g^{2}}\,\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}}+\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}.$ (25)
This form makes manifest the chiral structure of the graviweak theory and sets the stage for further unification via larger gauge groups, as we now show.
IV Extension of Lorentz Group – $\boldsymbol{SO(1,3+N)}$
Bearing this motivation in mind, in this Section we extend the Lorentz group to $SO(1,3+N)$. A major implication will be the appearance of possible interactions between the gravitational 2-form curvature $R^{a}_{~{}b}$ of $SO(1,3)$ and the Yang-Mills field strength $F^{ij}$ of a $SO(N)$ group. This is a remarkable novelty with regards to previous graviweak models.
IV.1 Structure of the Gauge Fields
Let us start by considering the Lie algebra decomposition:
$\mathfrak{so}(1,3+N)\cong\mathfrak{so}(1,3)\oplus\mathfrak{so}(N).$ (26)
The generators $M_{AB}$ obey the commutation relation $[M_{AB},M_{CD}]=\eta_{AD}M_{BC}-\eta_{AC}M_{BD}-\eta_{BD}M_{AC}+\eta_{BC}M_{AD}$ and have indices running as $A,B,C...=a,b,c,...,i,j,k,...=0,...,3+N$, where $a,b,c,..=0,1,2,3$ are in $\mathfrak{so}(1,3)$ and $i,j,k,...=4,...,3+N$ in $\mathfrak{so}(N)$. We therefore have $\frac{(N+4)(N+3)}{2}$ generators, which break into the 6 generators of $SO(1,3)$ plus the $\frac{N(N-1)}{2}$ generators of $SO(N)$.
The gauge field is valued in the Lie algebra of $\mathfrak{so}(1,3+N)$, with gauge field components $\mathcal{A}^{A}$ and its field strength is given by:
$\mathcal{F}=d\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=\frac{1}{2}\mathcal{F}_{AB}M^{AB},$
where $\mathcal{A}=\omega+A+\phi$, $\omega=\frac{1}{2}\omega_{ab}M^{ab}$ is the spin connection, $A=\frac{1}{2}A_{ij}M^{ij}$ is the Yang-Mills connection and $\phi=\phi_{ai}M^{ai}$ the coset field. We can write the total curvature in components as:
$\mathcal{F}_{AB}=d\mathcal{A}_{AB}+\mathcal{A}_{AC}\wedge\mathcal{A}^{C}_{~{}B},$ (27)
and more explicitly:
$\mathcal{F}^{a}_{~{}b}=R^{a}_{~{}b}+\phi^{a}_{~{}i}\wedge\phi^{i}_{~{}b},$ (28)
$\mathcal{F}^{i}_{~{}j}=F^{i}_{~{}j}+\phi^{i}_{~{}a}\wedge\phi^{a}_{~{}j},$ (29)
$\mathcal{F}^{a}_{~{}i}=d\phi^{a}_{~{}i}+\omega^{a}_{~{}c}\wedge\phi^{c}_{~{}i}+\phi^{a}_{~{}j}\wedge A^{j}_{~{}i}.$ (30)
The Killing form for this group is
[…]
With these tools in hand we now consider a specific realization focusing on how it augments the construction in Sections II and III.
IV.2 The $\mathbb{SO(1,6)}$ case
As an example let us consider the larger group $SO(1,6)$. The total curvature is defined as in (27) and an action quadratic in the curvature that we can consider in order to couple Yang-Mills and Einstein-Cartan is
$S$ $=\int\epsilon_{A_{1}...A_{7}}\mathcal{F}^{A_{1}A_{2}}\wedge\mathcal{F}^{A_{3}A_{4}}V^{A_{5}A_{6}A_{7}}$ (31)
$+\mathcal{F}^{AB}\wedge*\mathcal{F}_{AB}$
where we have introduced the auxiliary 0-form fields $V^{ABC}$. The second term corresponds to the kinetic term of the gauge field with the spacetime hodge dual; the first term can be written explicitly in terms of its $\mathfrak{so}(1,3)$ and $\mathfrak{so}(3)$ components as
$S_{\text{int}}$ $=S_{4,0}+S_{3,1}+S_{2,2}^{(a)}+S_{2,2}^{(b)}+S_{1,3},$ (32)
with
$S_{4,0}$ $=\int\;\epsilon_{abcd\,ijk}\;(\mathcal{F}^{ab}\wedge\mathcal{F}^{cd})\;V^{ijk},$ (34)
$S_{3,1}$ $=-12\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ab}\wedge\mathcal{F}^{ci}\bigr{)}\;V^{djk},$ (35)
$S_{2,2}^{(a)}$ $=6\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ab}\wedge\mathcal{F}^{ij}\bigr{)}\;V^{cd\,k},$ (36)
$S_{2,2}^{(b)}$ $=-12\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ai}\wedge\mathcal{F}^{bj}\bigr{)}\;V^{cd\,k},$ (37)
$S_{1,3}$ $=-4\int\;\epsilon_{abcd\,ijk}\;\bigl{(}\mathcal{F}^{ai}\wedge\mathcal{F}^{jk}\bigr{)}\;V^{bcd}.$ (38)
At this stage, one can consider breaking the symmetry in order to recover the Einstein-Cartan gravity sector along with the new interaction Einstein-Cartan/Yang-Mills interaction term. The first one arises from $S_{4,0}$ with the following symmetry breaking ansatz:
$\phi^{a}_{~{}i}=\frac{1}{\ell}e^{a}n_{i},\,\,\,\,n_{i}n^{i}=1,$ (39)
where $e^{a}$ is the $SO(1,3)$ 1-form tetrad, $n_{i}$ a normalized vector in $SO(3)$ and $\ell$ will be related to the cosmological constant $\Lambda$. Hence, the curvature terms become
$\mathcal{F}^{a}_{~{}b}=R^{a}_{~{}b}+\ell^{-2}e^{a}\wedge e_{b},$ (40)
$\mathcal{F}^{i}_{~{}j}=F^{i}_{~{}j},$ (41)
$\mathcal{F}^{a}_{~{}i}=\frac{1}{\ell}\Bigl{(}T^{a}\,n_{i}+e^{a}\wedge n_{j}A^{j}{}_{i}\Bigr{)},$ (42)
where the first equation has the same form as the one in MacDowell-Mansouri theories MacDowell:1977jt . We need an additional condition on $V^{ABC}$ in order to eliminate the remaining terms in the action. This can be achieved with:
$V^{ABC}=\varphi\epsilon^{ijk}\delta^{A}_{i}\delta^{B}_{j}\delta^{c}_{k}+\delta^{A}_{a}\delta^{B}_{b}\delta^{C}_{k}V^{abk}.$ (43)
Implementing the above ansatz for $\phi$ and $V$, action $S_{4,0}$ yields the Einstein-Cartan action plus the Euler term:
$S_{EC}$ $=6\int\varphi~{}\epsilon_{abcd}(R^{ab}+\ell^{-2}e^{a}\wedge e^{b})$ (44)
$\wedge(R^{cd}+\ell^{-2}e^{c}\wedge e^{d}),$
with the identifications $\Lambda=-3/\ell^{2}$ and $\varphi=\frac{1}{16\pi G}\frac{\ell^{2}}{24}$.
The second result of this broken symmetry is the Einstein-Cartan-Yang-Mills coupling that stems from action $S_{2,2}^{(a)}$ which takes the form:
$S_{RF}$ $=6\int\;\epsilon_{abcd}(R^{ab}+\ell^{-2}e^{a}\wedge e^{b})\wedge\epsilon_{ijk}F^{ij}V^{cd\,k}.$ (45)
The full action after symmetry breaking can be recasted as:
$S$ $=\frac{1}{16\pi G}\int e^{a}\wedge e^{b}\wedge\star R_{ab}-\frac{3}{2\Lambda}R^{ab}\wedge\star R_{ab}$ (46)
$-\frac{\Lambda}{12}\epsilon_{abcd}e^{a}\wedge e^{b}\wedge e^{c}\wedge e^{d}$
$+6\int\;\epsilon_{abcd}(R^{ab}-\frac{\Lambda}{3}e^{a}\wedge e^{b})\wedge\epsilon_{ijk}F^{ij}V^{cd\,k}$
$+\mathcal{F}^{AB}\wedge*\mathcal{F}_{AB}.$
and it describes graviweak theory. This Lagrangian possesses nonminimal couplings between the curvature form and the Yang-Mills fields Essig:2013lka .
V Conclusions and Outlook
In this paper we furthered the graviweak approach Alexander:0706.4481 ; Alexander:2012ge in two ways. First, we wedded it with the Higgs mechanism higgs , explaining how a graviweak construction could arise from a parity invariant action, with the symmetry broken by a Higgs field with a $Z_{2}$ symmetric potential. Second, we linked the ensuing construction with ideas inspired by MacDowell:1977jt , and investigated whether the $SU(2)$ employed in graviweak could be embedded in a larger group, which is then spontaneously broken. The first development allows for graviweak theory to be understood within a framework ubiquitous in unification theories: spontaneous symmetry breaking. The second is motivated mainly by the drive to seek experimental predictions which might prove or disprove this theory. We found that in most constructions arising from larger groups a non-minimal coupling between the curvature and the Yang-Mills fields is inevitable. At a quantum level this would imply a vertex involving the graviton and the weak gauge bosons.
Obviously this is but a proof of concept for testability and falsifiability. Still, one wonders if the possible production of a graviton background in the early Universe might not get modified by the presence of the weak sector. Parity violations in the graviton background would have dramatic effects on the CMB polarization Contaldi:2008iw ; Lue:1999rpc ; Saito:0705.3701 . With a new generation of gravitational wave detectors and CMB experiments expected to come online in the next decade, graviweak could be put to the test.
Further work could explore detailed phenomenological models to quantify precisely how graviweak-induced parity violations could manifest observationally. Additionally, theoretical investigations could seek a deeper understanding of the interplay between graviweak symmetry breaking and established cosmological models, particularly in inflationary and pre-inflationary epochs. Exploring connections with quantum gravity approaches such as loop quantum gravity Rovelli:2004tv ; Alexander:2011qfa might also provide valuable insights into the fundamental structure of space-time and particle interactions. Finally, computational simulations of the early universe that incorporate graviweak effects could greatly enhance our predictive power and help guide future experimental efforts.
Acknowledgements.
We thank Aaron Hui and Jacob Kuntzleman for helpful conversations. BA was supported by FCT Grant No. 2021.05694.BD and JM was partly supported by STFC Consolidated Grant ST/T000791/1.
Appendix A Translation into the two-spinor notation
Consider the self- and anti-self-dual decomposition of all objects as in equation (2). We want to write the gravity sector in terms of its self- and anti-self-dual parts in both representations. The area form is given in terms of the tetra by $\Sigma^{ab}=e^{a}\wedge e^{b}$. Therefore, we obtain for the Einstein-Cartan sector:
$\frac{1}{2}\epsilon_{abcd}\Sigma^{ab}\wedge R^{cd}=i(\Sigma^{(-)ab}\wedge R^{(-)}_{ab}-\Sigma^{(+)ab}\wedge R^{(+)}_{ab}).$ (47)
We use the Infeld–van der Waerden map, where spacetime indices $a=0,1,2,3$ are replaced by spinor indices via the soldering form $\sigma^{a}_{AA^{\prime}}$ and consider the following relations:
$\epsilon_{AA^{\prime}BB^{\prime}CC^{\prime}DD^{\prime}}=-i(\epsilon_{AC}\epsilon_{BD}\epsilon_{A^{\prime}D^{\prime}}\epsilon_{B^{\prime}C^{\prime}}-\epsilon_{AD}\epsilon_{BC}\epsilon_{A^{\prime}C^{\prime}}\epsilon_{B^{\prime}D^{\prime}})$ (48)
$\epsilon^{AB}\epsilon_{BC}=\epsilon^{A}_{~{}C}=-\delta^{A}_{C}$ (49)
$\epsilon^{AB}\epsilon_{CB}=\epsilon_{C}^{~{}A}=\delta^{A}_{C}.$ (50)
We also have the following decompositions, again in terms of self- and anti-self-dual components:
$R^{CC^{\prime}DD^{\prime}}=R^{(-)CC^{\prime}DD^{\prime}}+R^{(+)CC^{\prime}DD^{\prime}}=\epsilon^{C^{\prime}D^{\prime}}F^{CD}+\epsilon^{CD}\bar{F}^{C^{\prime}D^{\prime}}$ (51)
$\Sigma^{CC^{\prime}DD^{\prime}}=\Sigma^{(-)CC^{\prime}DD^{\prime}}+\Sigma^{(+)CC^{\prime}DD^{\prime}}=\epsilon^{C^{\prime}D^{\prime}}\Sigma^{CD}+\epsilon^{CD}\bar{\Sigma}^{C^{\prime}D^{\prime}},$ (52)
and we write $\Sigma^{CC^{\prime}DD^{\prime}}=e^{CC^{\prime}}\wedge e^{DD^{\prime}}$. The mapping introduced above should yield
$\frac{1}{2}\epsilon_{abcd}\Sigma^{ab}\wedge R^{cd}=\frac{1}{2}\epsilon_{AA^{\prime}BB^{\prime}CC^{\prime}DD^{\prime}}\Sigma^{AA^{\prime}BB^{\prime}}\wedge R^{CC^{\prime}DD^{\prime}},$ (53)
where the left hand side was computed above and the right hand side is, after several simplifications using the $\epsilon$ identities,
$\frac{1}{2}\epsilon_{AA^{\prime}BB^{\prime}CC^{\prime}DD^{\prime}}\Sigma^{AA^{\prime}BB^{\prime}}\wedge R^{CC^{\prime}DD^{\prime}}=2i(\Sigma^{AB}\wedge F_{AB}-\bar{\Sigma}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}}).$ (54)
Comparing this with (47) we arrive at
$i(\Sigma^{(-)ab}R_{(-)ab}-\Sigma^{(+)ab}R_{(+)ab})=2i(\Sigma^{AB}F_{AB}-\bar{\Sigma}^{A^{\prime}B^{\prime}}\bar{F}_{A^{\prime}B^{\prime}}).$ (55)
From (52) we get
$\Sigma^{AB}=-\frac{1}{2}e^{AC^{\prime}}\wedge e^{B}_{\,\,\,\,C^{\prime}}$ (56)
$\bar{\Sigma}^{A^{\prime}B^{\prime}}=-\frac{1}{2}e^{CA^{\prime}}\wedge e_{C}^{\,\,\,\,B^{\prime}}$ (57)
and so the SD part of the action can be written in terms of the tetrad components as
$\mathcal{L}_{\text{GR}}=\frac{i}{8\pi G}\,\Sigma^{AB}\wedge F_{AB}=-\frac{i}{16\pi G}\,e^{AA^{\prime}}\wedge e^{B}_{\,\,\,\,A^{\prime}}\wedge F_{AB}.$
The kinetic Yang-Mills term becomes:
$\mathcal{L}_{\text{YM}}=\frac{1}{4g_{+}^{2}}R^{(+)\,ab}\wedge\star R^{(+)}_{ab}=-\frac{i}{2g_{+}^{2}}\bar{F}^{A^{\prime}B^{\prime}}\wedge\bar{F}_{A^{\prime}B^{\prime}}.$ (58)
The fermionic Lagrangian can be written explicitly in the spinor indices as
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\star e^{a}\wedge\bar{\psi}_{L}\gamma_{a}D^{(+)}\psi_{L}=\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{A}\sigma^{a\,AA^{\prime}}D^{(+)}\psi_{A^{\prime}}$ (59)
$=\frac{i}{2}\,\star e^{AA^{\prime}}\wedge\bar{\psi}_{A}\,D^{(+)}\psi_{A^{\prime}}$ (60)
where
$D^{(+)}\psi_{A^{\prime}}=d\psi_{A^{\prime}}+\frac{1}{4}\bar{A}_{A^{\prime}}^{\,\,\,\,B^{\prime}}\psi_{B^{\prime}},$ (61)
$\bar{A}_{A^{\prime}}^{\,\,\,\,B^{\prime}}\psi_{B^{\prime}}=\omega^{(+)ab}\sigma_{a\,CA^{\prime}}\sigma_{b}^{CB^{\prime}}\psi_{B^{\prime}}$ (62)
and
$\star e_{a}\sigma^{a}_{AA^{\prime}}=\frac{1}{3!}\epsilon_{abcd}\wedge e^{b}\wedge e^{c}\wedge e^{d}~{}\sigma^{a}_{AA^{\prime}}=\frac{i}{3}e_{A}^{\,\,\,\,C^{\prime}}\wedge e_{DC^{\prime}}\wedge e^{D}_{\,\,\,\,A^{\prime}}\equiv\star e_{AA^{\prime}}.$